OpenStudy (anonymous):

Find two consecutive even integers such that smaller added to two times the larger gives the sum of 40.

OpenStudy (anonymous):

10 and 15

OpenStudy (anonymous):

my guess is 10 and 12 10 + 2*12 = 10 + 24 = 34 so my guess is wrong

OpenStudy (anonymous):

10+30=40

OpenStudy (anonymous):

try 12 and 14

OpenStudy (anonymous):

15x2=30+10=40

OpenStudy (anonymous):

two consecutive integers can be x and x+1 let x+1 be the smaller number then 2x+(x+1) = 40 solve it and get your answers

OpenStudy (anonymous):

even even integers!!

OpenStudy (anonymous):

two consecutive EVEN intgegers

OpenStudy (anonymous):

you were right @satellite73

OpenStudy (anonymous):

sorry two consecutive even integers would be x and x+2

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

lol yeah i know

OpenStudy (anonymous):

just guess with a problem like this, forget algebra

OpenStudy (anonymous):

ok,i didn't read it all the way sorry, @sattelite73,you were correct.